OpenStudy (anonymous):

Factor into a product of binomial X^8-9

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

i think this is what they are asking:\[a^2 - b^2 = (a+b)(a - b)\] a^2 = x^8 b^2 = 9 a = x^4 b = 3\[x^8 - 9 = (x^4 +3)(x^4 - 3)\]

OpenStudy (anonymous):

Thank you this was very helpful